OpenStudy (anonymous):

find the laplace of f(t) = e^(-2t) * sin(3pi*t)

OpenStudy (anonymous):

integral from 0 to infinity?

OpenStudy (anonymous):

okay so F(s +2)

OpenStudy (anonymous):

so F(s) + F(2)

OpenStudy (anonymous):

how do i find F(s+2)

OpenStudy (anonymous):

okay and whats the L of coskt?

OpenStudy (anonymous):

Wait!!! Is it convolution???

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

apply the translation theorem

OpenStudy (anonymous):

L(f*g) = F(s)G(s) :S ^Convolution

hartnn (hartnn):

L[f1(t)*f2(t)] = L [f1(t)] L [f2(t)] so just find L[e^{-2t}] and L[sin(3pi*t)] and multiply them